Probability Rules
Basic probability rules:
Rule 1: For any event A: 0≤P(A)≤1
Rule 2 (Sample Space Rule): P(S)=1 — some outcome always occurs
Rule 3 (Complement Rule):
P(Ac)=1−P(A)
where
Ac is the complement of
A (event A does NOT occur)
Rule 4 (Addition Rule for Mutually Exclusive Events):
If
A and
B cannot both occur:
P(A or B)=P(A)+P(B) Equally likely outcomes
P(A)=total number of outcomes in Snumber of outcomes in A Example: Roll a fair die: P(even)=63=0.5